# VoiceTextBlender: Augmenting Large Language Models with Speech Capabilities via Single-Stage Joint Speech-Text Supervised Fine-Tuning

### v1

@@ Abstract
Recent studies have augmented large language models (LLMs) with speech capabilities, leading to the development of speech language models (SpeechLMs). Earlier SpeechLMs focused on single-turn speech-based question answering (QA), where user input comprised a speech context and a text question. More recent studies have extended this to multi-turn conversations, though they often require complex, multi-stage supervised fine-tuning (SFT) with diverse data. Another critical challenge with SpeechLMs is catastrophic forgetting—where models optimized for speech tasks suffer significant degradation in text-only performance. To mitigate these issues, we propose a novel single-stage joint speech-text SFT approach on the low-rank adaptation (LoRA) of the LLM backbone. Our joint SFT combines text-only SFT data with three types of speech-related data: speech recognition and translation, speech-based QA, and mixed-modal SFT. Compared to previous SpeechLMs with 7B or 13B parameters, our 3B model demonstrates superior performance across various speech benchmarks while preserving the original capabilities on text-only tasks. Furthermore, our model shows emergent abilities of effectively handling previously unseen prompts and tasks, including multi-turn, mixed-modal inputs. [^1] Our scripts for data generation, codebase for training and inference, and pre-trained model weights will be publicly released upon acceptance.
@@
Yifan Peng1thanks: Equal contribution. Work done while Yifan was an intern at NVIDIA., Krishna C. Puvvada2, Zhehuai Chen2, Piotr Zelasko2, He Huang2, Kunal Dhawan2, Ke Hu2, Shinji Watanabe1, Jagadeesh Balam2, Boris Ginsburg2 1Carnegie Mellon University, 2NVIDIA
@@ 1 Introduction
  • We propose a single-stage joint speech-text SFT strategy for training SpeechLMs, which simplifies the training process, preserves the LM’s original text-only performance, and delivers strong results on speech tasks. Our 3B model outperforms previous 7B or 13B SpeechLMs on most evaluated benchmarks.
@@ 2 Related Work
SpeechLM training. SpeechLMs are typically trained in multiple stages with supervised data. Gong et al. find that an appropriate curriculum is crucial to train their LTU models. Specifically, they propose a four-stage training procedure that gradually increases the number of learnable parameters and the complexity of the training tasks. Tang et al. adopt a three-stage training pipeline for their SALMONN models: pre-training, instruction tuning, and activation tuning. SpeechVerse Das et al. (2024) observes that training all learnable parameters from scratch on diverse speech tasks often leads to divergence. Hence, they propose two-stage training with gradually increased learnable modules and speech tasks. While multi-stage curriculum learning methods are intuitive and enhance training stability, they significantly increase the complexity of design choices. These approaches require extensive tuning and heuristics to determine the appropriate order for updating modules and assigning tasks. In this work, we adopt a single-stage training strategy that combines text-only SFT data with various mixed-modal SFT data. Our approach streamlines the training pipeline while achieving strong performance across diverse benchmarks.
Catastrophic forgetting in SpeechLM. Most SpeechLMs are optimized for speech tasks, often at the expense of their original text capabilities, a phenomenon known as catastrophic forgetting. To preserve the text-only performance of instruction-tuned LMs, some studies freeze the backbone LM Wang et al. (2023a); Fathullah et al. (2024); Dubey et al. (2024). However, as discussed in Section 4.3 and by Wang et al., this approach may degrade performance on speech tasks. Many other works use parameter-efficient fine-tuning methods, such as LoRA adapters Hu et al. (2022), in an attempt to mitigate catastrophic forgetting Gong et al. (2024); Das et al. (2024). However, our experiments in Section 4.3 reveal that merging the LoRA parameters into the original model parameters significantly degrades performance on text-only benchmarks, demonstrating that LoRA alone does not ensure preservation of the model’s original capabilities. Inspired by recent work in vision language models, such as VILA Lin et al. (2024), we propose a joint speech-text SFT approach, which integrates text-only SFT data with speech-related SFT data. Our method preserves text-only performance while achieving strong results on newly added speech tasks.
@@ 3.2.2 Speech-based QA from ASR Data
To enable general QA capabilities about speech, we create speech-based QA data from English ASR data by prompting a pre-trained LLM. As shown in Figure 3(a), we provide the transcript to an LLM [^2] We use the official chat version of Gemma2 27B: https://huggingface.co/google/gemma-2-27b-it and prompt it to generate a question-answer pair based on the provided context. During training, the text question and corresponding speech context are given as input, while the model is trained to predict the text answer. This approach has been used in prior studies Tang et al. (2024); Gong et al. (2024); Noroozi et al. (2024), but we scale it up to 20k hours of audio.
@@ 3.2.3 Mixed-Modal SFT Data with TTS
This data has more flexible input formats than SQA, making speech and text inputs interchangeable. Instructions can now be conveyed through speech rather than being limited to text. Additionally, text SFT data typically covers more diverse instructions, and the responses maintain high-quality language and style. This contributes to the overall quality of our generated mixed-modal data. However, a potential limitation of this data type is that the audio is synthetic, reflecting only the limited acoustic conditions provided by the TTS model.
@@ 5 Conclusion
Due to license issues, some of the training data cannot be directly released. Instead, we provide statistics about those data and describe the details about our training procedure. For SQA and mixed-modal SFT, we plan to release the data generation scripts.

### v2

@@ Abstract
Recent studies have augmented large language models (LLMs) with speech capabilities, leading to the development of speech language models (SpeechLMs). Earlier SpeechLMs focused on single-turn speech-based question answering (QA), where user input comprised a speech context and a text question. More recent studies have extended this to multi-turn conversations, though they often require complex, multi-stage supervised fine-tuning (SFT) with diverse data. Another critical challenge with SpeechLMs is catastrophic forgetting, where models optimized for speech tasks suffer significant degradation in text-only performance. To mitigate these issues, we propose a novel single-stage joint speech-text SFT approach on the low-rank adaptation (LoRA) of the LLM backbone. Our joint SFT combines text-only SFT data with three types of speech-related data: speech recognition and translation, speech-based QA, and mixed-modal SFT. Compared to previous SpeechLMs with 7B or 13B parameters, our 3B model demonstrates superior performance across various speech benchmarks while preserving the original capabilities on text-only tasks. Furthermore, our model shows emergent abilities of effectively handling previously unseen prompts and tasks, including multi-turn, mixed-modal inputs. [^1] We will publicly release our data generation scripts, training code, and pre-trained model weights: https://github.com/pyf98/NeMo_VoiceTextBlender
@@
Yifan Peng1thanks: Equal contribution. Work done while Yifan was an intern at NVIDIA., Krishna C. Puvvada2, Zhehuai Chen2, Piotr Zelasko2, He Huang2, Kunal Dhawan2, Ke Hu2, Shinji Watanabe1, Jagadeesh Balam2, Boris Ginsburg2 1Carnegie Mellon University, 2NVIDIA Correspondence: pengyf21@gmail.com, kpuvvada@nvidia.com, zhehuaic@nvidia.com
@@ 1 Introduction
  • We propose a single-stage joint speech-text SFT strategy for training SpeechLMs, which simplifies the training process, preserves the LM’s original text-only performance, and delivers strong results on speech tasks. [^2] The current work mainly focuses on the linguistic content of human speech. In the future, we will expand to diverse attributes such as speaker identity Wu et al. (2024). Our 3B model outperforms previous 7B or 13B SpeechLMs on most evaluated benchmarks.
@@ 2 Related Work
SpeechLM training. SpeechLMs are typically trained in multiple stages with supervised data. Gong et al. find that an appropriate curriculum is crucial to train their LTU models. Specifically, they propose a four-stage training procedure that gradually increases the number of learnable parameters and the complexity of the training tasks. Tang et al. adopt a three-stage training pipeline for their SALMONN models: pre-training, instruction tuning, and activation tuning. SpeechVerse Das et al. (2024) observes that training all learnable parameters from scratch on diverse speech tasks often leads to divergence. Hence, they propose two-stage training with gradually increased learnable modules and speech tasks. While multi-stage curriculum learning methods are intuitive and enhance training stability, they significantly increase the complexity of design choices. These approaches require extensive tuning and heuristics to determine the appropriate order for updating modules and assigning tasks. In this work, we adopt a single-stage training strategy that combines text-only SFT data with various mixed-modal SFT data. Our approach streamlines the training pipeline while achieving strong performance across diverse benchmarks. [^3] A recent work, AudioChatLlama Fathullah et al. (2024), also conducts single-stage training. However, it is not explicitly trained on diverse speech tasks. Instead, it uses ASR data and relies on the modal-invariance trick, which differs greatly from our training objective.
Catastrophic forgetting in SpeechLM. Most SpeechLMs are optimized for speech tasks, often at the expense of their original text capabilities, a phenomenon known as catastrophic forgetting. To preserve the text-only performance of instruction-tuned LMs, some studies freeze the backbone LM Wang et al. (2023a); Fathullah et al. (2024); Dubey et al. (2024). However, as discussed in Section 4.3 and by Wang et al., this approach may degrade performance on speech tasks. Many other works use parameter-efficient fine-tuning methods, such as LoRA adapters Hu et al. (2022), to mitigate catastrophic forgetting Gong et al. (2024); Das et al. (2024). However, our experiments in Section 4.3 reveal that merging the LoRA parameters into the original model parameters significantly degrades performance on text-only benchmarks, demonstrating that LoRA alone does not ensure preservation of the model’s original capabilities. A recent study in vision language models, VILA Lin et al. (2024), finds that incorporating text SFT data in their multi-stage training paradigm mitigates catastrophic forgetting. However, it only considers the vision modality, not speech. Inspired by this line of work, we propose a joint speech-text SFT approach, which integrates text-only SFT data with speech-related SFT data. Our method preserves text-only performance while achieving strong results on newly added speech tasks.
@@ 3.2.2 Speech-based QA from ASR Data
To enable general QA capabilities about speech, we create speech-based QA data from English ASR data by prompting a pre-trained LLM. As shown in Figure 3(a), we provide the transcript to an LLM [^4] https://huggingface.co/google/gemma-2-27b-it and prompt it to generate a question-answer pair based on the provided context. During training, the text question and corresponding speech context are given as input, while the model is trained to predict the text answer. This approach has been used in prior studies Tang et al. (2024); Gong et al. (2024); Noroozi et al. (2024), but we scale it up to 20k hours of audio.
@@ 3.2.3 Mixed-Modal SFT Data with TTS
This data has more flexible input formats than SQA, making speech and text inputs interchangeable. Instructions can now be conveyed through speech rather than being limited to text. Additionally, text SFT data typically covers more diverse instructions, and the responses maintain high-quality language and style. This contributes to the overall quality of our generated mixed-modal data. However, a potential limitation of this data type is that the audio is synthetic, reflecting only the limited acoustic conditions provided by the TTS model. [^5] For simplicity, this work uses a single TTS model. Future work can explore multiple TTS models with diverse speakers and emotions to enhance robustness.
@@ 5 Conclusion
The text pre-training data of Gemma is not publicly released, which might raise concerns. Due to license issues, some of the speech training data cannot be directly released. Instead, we provide statistics about those data and describe the details about our training procedure. For SQA and mixed-modal SFT, we plan to release the data generation scripts.